# Analytic model of stable shock-like structures in laser interaction with underdense plasma for identifying of phase and polarization dependent regime of laser wakefield accelerators

### v1

@@ I Introduction
Laser Wakefield Accelerators (LWFA) can sustain extremely high accelerating gradients, providing a pathway to ultra-compact electron accelerators.lwfa0 ; lwfa1 ; lwfa2 ; lwfa3 Plasma electrons are pushed out by the radiation pressure of the driving ultra-intense laser, forming a region devoid of electrons called "plasma bubble" propagating with a speed close to speed of light.bubble1 ; bubble2 ; bubble3 ; bubble4 The acceleration electric field in "bubble" region can reach dozens of GeV$/m$.gev1 ; gev2
While the phase-averaged (ponderomotive) description of the plasma response to the laser is usually sufficientpond1 , phase and polarization dependent Carrier-Envelope Phase (CEP) effect can manifest when the laser self-steepens, increasing to peak amplitude over the time-scale of a single laser cycle.cep1 ; cep2 Under such circumstance, the plasma bubble undergoes asymmetric phase and polarization dependent periodic undulation.phase1 ; phase2 ; phase3 This enables control of electron injection and acceleration process giving rise to ultra-high current spatiotemporally structured electron beams. phasehigh1 ; phase1
@@
In this paper, we identify the existence of stable shock-like structure using an 1D Quasistatic code Wake1D under a wide range of laser-plasma parameters pic1 . Based on this finding, we derive an analytical expression that in turn describe the significance of CEP effect for a given laser-plasma parameter. The paper is organized as follows: in section II, we develop a theoretical model for a stable shock-like laser structure based on quasistatic model of laser-plasma interaction quasi1 ; former2 based on observations from a quasi-static particle-in-cell simulation WAKE1D. In section III, we analyze the steepening parameters using normalized expressions from this model, which is in turn used to parametrize the importance of CEP for a given laser-plasma parameters.
@@ II Persistence of a stable shock front structure
$\left(ik_{0}\frac{2\partial}{\partial t}-\frac{2\partial^{2}}{\partial t\partial\xi}\right)\hat{A}_{\perp}=\chi\hat{A}_{\perp}$ (1)
,where the $\chi=\frac{1}{\psi+1}$ , represents the plasma density accounting for the relativistic transparency .density1
@@
$\frac{\partial^{2}\psi}{\partial\xi^{2}}=\frac{1+|\hat{A}_{\perp}|^{2}}{2(1+\psi)^{2}}-\frac{1}{2}$ (2)
Inserting $\hat{A}_{\perp}=a\exp(i\theta)$, where $a$ and $\theta$ are both real numbers, into Eq(2):
$\left(ik_{0}\frac{2\partial}{\partial t}-\frac{2\partial^{2}}{\partial t\partial\xi}\right)(a\exp(i\theta))=\frac{1}{1+\psi}(a\exp(i\theta))$ (3)
Our primary focus is identification of a stable shock-like structure propagating at a constant sub-luminal velocity. We define the speed of front etching in the copropagating frame as $v_{etch}=1-v_{front}$, where $v_{front}$ represents the speed of the front structure in the laboratory frame. Taking note of the etching velocity found in the 1D simulation, we assume that the vector potential can be written as a function of a single variable $\xi-v_{etch}t$, resulting in the transformation of time and space: $\partial_{t}a=-v_{etch}\partial_{\xi}a$ and $\partial_{t}\theta=-v_{etch}\partial_{\xi}\theta$. We then convert Eq(3) into a function with only one independent variable $\xi$:
@@
$+a(\partial_{\xi}\theta^{2}-k_{0}\partial_{\xi}\theta-\partial_{\xi}^{2}a)=-\frac{1}{2v_{etch}}\frac{1}{1+\psi}a$ (5)
@@
$\begin{cases}\frac{\partial^{2}a}{\partial\xi^{2}}=(-\frac{k_{0}^{2}}{4}+\frac{1}{2v_{etch}}\frac{1}{1+\psi})a\\ \frac{\partial^{2}\psi}{\partial\xi^{2}}=\frac{a^{2}}{2(1+\psi)^{2}}+\frac{1}{2}(\frac{1}{(1+\psi)^{2}}-1)\\ \theta=\frac{k_{0}}{2}(\xi-v_{etch}t)\end{cases}$ (6)
@@
The accuracy of the assumptions we made can be evaluated using the WAKE1D PIC. We note that the configuration is determined by $a_{0}$ and $k_{0}$ in PIC simulation while it depends on $a_{0}$ and $v_{etch}$ in our analytic approach. The two different approaches describing the same structure means that there are only 2 degrees of freedom among the three parameters $a_{0}$, $k_{0}$ and $v_{etch}$. Thus, the relation of them can be a index of accuracy ; indeed, reasonable accuracy is found for a wide range of parameters for PIC and our approach. [Fig 1]
@@ III.1 Parameter-free theory of steepened shock-front
$\tilde{B}_{y}^{L}=\frac{\partial\tilde{A}_{x}}{\partial z}-\frac{\partial\tilde{A}_{z}}{\partial x}\approx\left(-\frac{\partial\hat{A}_{x}}{\partial\xi}+ik_{0}\hat{A}_{x}\right)\exp\left(-ik_{0}\xi\right)$ (7)
@@
$|\tilde{B}_{y}^{L}|^{2}=\frac{k_{0}^{2}}{4}a_{m}^{2}=k_{0}^{2}a_{0}^{2}$ (8)
, showing that the vector potential amplitude behind the shock converges to $a_{m}=2a_{0}$ and that the relation of $a_{0}$, $k_{0}$ and $v_{etch}$ can also be written as $a_{0}=\frac{a_{m}(k_{0},v_{etch})}{2}$.
From Eq (6), only periodical solutions exist when $\frac{1}{2v_{etch}}\geq\frac{k_{0}^{2}}{4}$ while steepening front solutions exists for opposite case $\frac{1}{2v_{etch}}<\frac{k_{0}^{2}}{4}$. This means that when steepening occurs, there is a upper limit of the etching velocity $v_{m}=\frac{2}{k_{0}^{2}}$. We define that $v_{etch}=(1-\Gamma)v_{m}$, where $\Gamma$ is a dimensionless quantity to quantifying the deviation from maximum etching velocity.
One way to understand this dimensionless quantity $\Gamma$ is using energy balance. The energy depletion rate of laser pulse can be estimated by the etching velocity; namely the amount of energy loss $W$ to the plasma at this shock-like front can be estimated by:energyloss1
$W=v_{etch}\frac{k_{0}^{2}a_{0}^{2}}{4\pi}$ (9)
@@
Due to the difficulty in obtaining the analytical solution, numerical approaches will be employed to derive the expressions. Extra coefficient in the equation set can be simplified by variable substitution. Assume the normalization units are $\hat{a}=\frac{\Gamma}{\sqrt{2(1-\Gamma)}}k_{0}$, $\hat{\xi}=2\sqrt{\frac{1-\Gamma}{\Gamma}}\frac{1}{k_{0}}$. And the normalized variables: $u=a/\hat{a}$, $f=(1-\frac{1}{1+\psi})/\Gamma$ and $x=\xi/\hat{\xi}$. Using these variables, Eq (6) can be recast to:
$\begin{cases}\frac{\partial^{2}u}{\partial x^{2}}=(1-f)u\\ \frac{\partial^{2}(1/(\Gamma^{2}-\Gamma f))}{\partial x^{2}}=u^{2}(1-\Gamma f)^{2}\end{cases}$ (10)
@@
$\frac{a_{m}}{\hat{a}}=u_{m}=u_{m}(\Gamma)$ (11)
,which can be obtained by one variable fitting.
$u_{m}=\frac{2}{\frac{1.19}{\Gamma}-1.46}\frac{\sqrt{2(1-\Gamma)}}{\Gamma}$ (12)
@@ III.2 Strength of CEP effect using the parameter-free expression
$u_{h}=u_{h}(\Gamma)$ (14)
$u_{x}=u_{x}(\Gamma)$ (15)
$\Delta_{u}=u_{h}/u_{x}=\Delta_{u}(\Gamma)$ (16)
, which can be obtained by fitting.
$\Delta_{u}=\frac{3.6}{\Gamma^{0.51}}\sqrt{\frac{\Gamma}{2(1-\Gamma)}}$ (18)
@@
$=\frac{3.56}{\Gamma^{0.5}}\frac{1}{k_{0}}$
@@
The result can be verified by the ratio of $k_{0}\Delta_{theory}$ gained by Eq (21) and $k_{0}\Delta_{calculation}$ obtained directly by computing of Eq (7) minus one. The error $|\frac{k_{0}\Delta_{theory}}{k_{0}\Delta_{calculation}}-1|$ is less than $2\%$. Utilising a more advanced fitting primitive can enhance the accuracy but will complicate the formula form.
@@
$k_{0}\Delta$ $=\frac{3.56}{\Gamma^{0.5}}$ (20)
$\frac{k_{0}}{a_{0}}$ $=\frac{1.19}{\Gamma}-1.46$ (21)
@@
$k_{0}\Delta=3.26(\frac{k_{0}}{a_{0}}+1.46)^{0.5}$ (22)
As a result, the contour line of $k_{0}\Delta$ is a straight line in $a_{0}$-$k_{0}$ space, which can be observed in Fig (3). Thus the threshold of CEP effect can be expressed as $a_{0}<k_{0}/2.25$.
In addition, the transverse momenta observed during the CEP effect can serve as another significant indicator: $\Lambda\propto\frac{a_{0}^{3}}{\sigma_{x}^{2}}\frac{1}{\sigma_{z}^{2}}\cos\left(\varphi+\varphi_{0}\right)$, where $\varphi$ is the CEP and $\varphi_{0}$ is the CEP offset, $\sigma_{x}$ and $\sigma_{z}$ are the transversal and longitudinal spot sizes of a Gaussian Pulse.cep1 While $\cos(\varphi+\varphi_{0})$ illustrates the dependence of electron injection on phase, the ratio $\frac{a_{0}^{3}}{\sigma_{x}^{2}}\frac{1}{\sigma_{z}^{2}}$ quantifies the strength of the coupling between injection and phase. To evaluate the strength in one-dimensional scenario, the term $a_{0}^{3}/(k_{0}\Delta)^{2}$ in $a_{0}-k_{0}$ space is also plotted in Fig (3).

### v2

@@ I Introduction
Laser Wakefield Accelerators (LWFA) can sustain extremely high accelerating gradients, providing a pathway to ultra-compact electron accelerators lwfa0 ; lwfa1 ; lwfa2 ; lwfa3 . Plasma electrons are pushed out by the radiation pressure of the driving ultra-intense laser, forming a region devoid of electrons called "plasma bubble" propagating with a speed close to speed of light bubble1 ; bubble2 ; bubble3 ; bubble4 . The acceleration electric field in "bubble" region can reach dozens of GeV$/m$ gev1 ; gev2 .
While the phase-averaged (ponderomotive) description of the plasma response to the laser is usually sufficient pond1 , phase and polarization dependent Carrier-Envelope Phase (CEP) effect can manifest when the laser self-steepens, increasing to peak amplitude over the time-scale of a single laser cycle cep1 ; cep2 . Under such circumstance, the plasma bubble undergoes asymmetric phase and polarization dependent periodic undulation phase1 ; phase2 ; phase3 . This enables control of electron injection and acceleration process giving rise to ultra-high current spatiotemporally structured electron beams phasehigh1 ; phase1 .
@@
In this paper, we identify the existence of stable shock-like structure using an 1D Quasistatic code Wake1D under a wide range of laser-plasma parameters pic1 . Based on this finding, we derive an analytical expression that in turn describe the significance of CEP effect for a given laser-plasma parameter. The paper is organized as follows: in section II, we develop a theoretical model for a stable shock-like laser structure based on quasistatic model of laser-plasma interaction quasi1 ; former2 , and observations from a quasi-static particle-in-cell simulation WAKE1D. In section III, we analyze the steepening parameters using normalized expressions from this model, which is in turn used to parametrize the importance of CEP for a given laser-plasma parameters.
@@ II Persistence of a stable shock front structure
$\left(ik_{0}\frac{2\partial}{\partial t}-\frac{2\partial^{2}}{\partial t\partial\xi}\right)\hat{A}_{\perp}=\chi\hat{A}_{\perp},$ (1)
where the $\chi=\frac{1}{\psi+1}$ , represents the plasma density accounting for the relativistic transparency density1 .
@@
$\frac{\partial^{2}\psi}{\partial\xi^{2}}=\frac{1+|\hat{A}_{\perp}|^{2}}{2(1+\psi)^{2}}-\frac{1}{2}.$ (2)
Inserting $\hat{A}_{\perp}=a\exp(i\theta)$, where $a$ and $\theta$ are both real numbers, into Eq. (2):
$\left(ik_{0}\frac{2\partial}{\partial t}-\frac{2\partial^{2}}{\partial t\partial\xi}\right)(a\exp(i\theta))=\frac{1}{1+\psi}(a\exp(i\theta)).$ (3)
Our primary focus is identification of a stable shock-like structure propagating at a constant sub-luminal velocity. We define the speed of front etching in the copropagating frame as $v_{etch}=1-v_{front}$, where $v_{front}$ represents the speed of the front structure in the laboratory frame. Taking note of the etching velocity found in the 1D simulation, we assume that the vector potential can be written as a function of a single variable $\xi-v_{etch}t$, resulting in the transformation of time and space: $\partial_{t}a=-v_{etch}\partial_{\xi}a$ and $\partial_{t}\theta=-v_{etch}\partial_{\xi}\theta$. We then convert Eq. (3) into a function with only one independent variable $\xi$:
@@
$+a(\partial_{\xi}\theta^{2}-k_{0}\partial_{\xi}\theta-\partial_{\xi}^{2}a)=-\frac{1}{2v_{etch}}\frac{1}{1+\psi}a.$ (5)
@@
$\begin{cases}\frac{\partial^{2}a}{\partial\xi^{2}}=(-\frac{k_{0}^{2}}{4}+\frac{1}{2v_{etch}}\frac{1}{1+\psi})a,\\ \frac{\partial^{2}\psi}{\partial\xi^{2}}=\frac{a^{2}}{2(1+\psi)^{2}}+\frac{1}{2}(\frac{1}{(1+\psi)^{2}}-1),\\ \theta=\frac{k_{0}}{2}(\xi-v_{etch}t).\end{cases}$ (6)
@@
The accuracy of the assumptions we made can be evaluated using the WAKE1D PIC. We note that the configuration is determined by $a_{0}$ and $k_{0}$ in PIC simulation while it depends on $a_{0}$ and $v_{etch}$ in our analytic approach. The two different approaches describing the same structure means that there are only 2 degrees of freedom among the three parameters $a_{0}$, $k_{0}$ and $v_{etch}$. Thus, the relation of them can be a index of accuracy ; indeed, reasonable accuracy is found for a wide range of parameters for PIC and our approach [Fig 1].
@@ III.1 Parameter-free theory of steepened shock-front
$\tilde{B}_{y}^{L}=\frac{\partial\tilde{A}_{x}}{\partial z}-\frac{\partial\tilde{A}_{z}}{\partial x}\approx\left(-\frac{\partial\hat{A}_{x}}{\partial\xi}+ik_{0}\hat{A}_{x}\right)\exp\left(-ik_{0}\xi\right).$ (7)
@@
$|\tilde{B}_{y}^{L}|^{2}=\frac{k_{0}^{2}}{4}a_{m}^{2}=k_{0}^{2}a_{0}^{2},$ (8)
showing that the vector potential amplitude behind the shock converges to $a_{m}=2a_{0}$ and that the relation of $a_{0}$, $k_{0}$ and $v_{etch}$ can also be written as $a_{0}=\frac{a_{m}(k_{0},v_{etch})}{2}$.
From Eq. (6), only periodical solutions exist when $\frac{1}{2v_{etch}}\geq\frac{k_{0}^{2}}{4}$ while steepening front solutions exists for opposite case $\frac{1}{2v_{etch}}<\frac{k_{0}^{2}}{4}$. This means that when steepening occurs, there is a upper limit of the etching velocity $v_{m}=\frac{2}{k_{0}^{2}}$. We define that $v_{etch}=(1-\Gamma)v_{m}$, where $\Gamma$ is a dimensionless quantity to quantifying the deviation from maximum etching velocity.
One way to understand this dimensionless quantity $\Gamma$ is using energy balance. The energy depletion rate of laser pulse can be estimated by the etching velocity; namely the amount of energy loss $W$ to the plasma at this shock-like front can be estimated by energyloss1 :
$W=v_{etch}\frac{k_{0}^{2}a_{0}^{2}}{4\pi}.$ (9)
@@
Due to the difficulty in obtaining the analytical solution, numerical approaches will be employed to derive the expressions. Extra coefficient in the equation set can be simplified by variable substitution. Assume the normalization units are $\hat{a}=\frac{\Gamma}{\sqrt{2(1-\Gamma)}}k_{0}$, $\hat{\xi}=2\sqrt{\frac{1-\Gamma}{\Gamma}}\frac{1}{k_{0}}$. And the normalized variables: $u=a/\hat{a}$, $f=(1-\frac{1}{1+\psi})/\Gamma$ and $x=\xi/\hat{\xi}$. Using these variables, Eq. (6) can be recast to:
$\begin{cases}\frac{\partial^{2}u}{\partial x^{2}}=(1-f)u,\\ \frac{\partial^{2}(1/(\Gamma^{2}-\Gamma f))}{\partial x^{2}}=u^{2}(1-\Gamma f)^{2}.\end{cases}$ (10)
@@
$\frac{a_{m}}{\hat{a}}=u_{m}=u_{m}(\Gamma),$ (11)
which can be obtained by one variable fitting.
$u_{m}=\frac{2}{\frac{1.19}{\Gamma}-1.46}\frac{\sqrt{2(1-\Gamma)}}{\Gamma}.$ (12)
@@ III.2 Strength of CEP effect using the parameter-free expression
$u_{h}=u_{h}(\Gamma),$ (14)
$u_{x}=u_{x}(\Gamma),$ (15)
$\Delta_{u}=u_{h}/u_{x}=\Delta_{u}(\Gamma),$ (16)
which can be obtained by fitting.
$\Delta_{u}=\frac{3.6}{\Gamma^{0.51}}\sqrt{\frac{\Gamma}{2(1-\Gamma)}}.$ (18)
@@
$=\frac{3.56}{\Gamma^{0.5}}\frac{1}{k_{0}}.$
@@
The result can be verified by the ratio of $k_{0}\Delta_{theory}$ gained by Eq. (21) and $k_{0}\Delta_{calculation}$ obtained directly by computing of Eq. (7) minus one. The error $|\frac{k_{0}\Delta_{theory}}{k_{0}\Delta_{calculation}}-1|$ is less than $2\%$. Utilising a more advanced fitting primitive can enhance the accuracy but will complicate the formula form.
@@
$k_{0}\Delta$ $=\frac{3.56}{\Gamma^{0.5}},$ (20)
$\frac{k_{0}}{a_{0}}$ $=\frac{1.19}{\Gamma}-1.46.$ (21)
@@
$k_{0}\Delta=3.26(\frac{k_{0}}{a_{0}}+1.46)^{0.5}.$ (22)
As a result, the contour line of $k_{0}\Delta$ is a straight line in $a_{0}$-$k_{0}$ space, which can be observed in Fig (3). Thus the threshold of CEP effect can be expressed as $a_{0}<k_{0}/2.25$. Although the exact form of the expression Eq. (22) is typically obtained through fitting for practical applications, the functional dependence of $k_{0}\Delta$ on $\frac{a_{0}}{k_{0}}$ is mathematically valid, providing a scaling relation.
In addition, the transverse momenta observed during the CEP effect can serve as another significant indicator: $\Lambda\propto\frac{a_{0}^{3}}{\sigma_{x}^{2}}\frac{1}{\sigma_{z}^{2}}\cos\left(\varphi+\varphi_{0}\right)$, where $\varphi$ is the CEP and $\varphi_{0}$ is the CEP offset, $\sigma_{x}$ and $\sigma_{z}$ are the transversal and longitudinal spot sizes of a Gaussian Pulse cep1 . While $\cos(\varphi+\varphi_{0})$ illustrates the dependence of electron injection on phase, the ratio $\frac{a_{0}^{3}}{\sigma_{x}^{2}}\frac{1}{\sigma_{z}^{2}}$ quantifies the strength of the coupling between injection and phase. To evaluate the strength in one-dimensional scenario, the term $a_{0}^{3}/(k_{0}\Delta)^{2}$ in $a_{0}-k_{0}$ space is also plotted in Fig (3).